American Food Services, Inc. leased a packaging machine from Barton and Barton Corporation. Barton and Barton completed constReg 1 Req 2 Req 3 and 4 Prepare the appropriate entries related to the lease on December 31, 2021 and 2023. (If no entry is rJournal entry worksheet Record the amortization of right-of-use asset for American Food Services. Note: Enter debits before cJournal entry worksheet < 1 2 3 4 Record the lease payment and interest expense for American Food Services. Note: Enter debitJournal entry worksheet < 1 2 3 4 Record the amortization of right-of-use asset for American Food Services. Note: Enter debit
Answer #1

B A 1 1) 2 Date 3 Jan 1,2021 Debit General Journal Right of use asset Lease Payable (To record the right of use asset and lea24 25 Date 26 Dec 31, 2021 27 28 Credit 4,00,000 8,61,881 12,61,881 General Journal Debit Interest Expense Lease Payable CashB D А 1 1) 2 Date 3 Jan 1,2021 Credit Debit 4000000 4. General Journal Right of use asset Lease Payable (To record the rightCredit Debit =C18 =D18 =C26+C27 29 30 =4000000/4 =C30 24 25 Date General Journal 26 Dec 31, 2021 Interest Expense 27 Lease Pa